### v2

# Microscopic model for aging of biocondensates

Abstract
Biomolecular condensates are membraneless compartments in the cell that are involved in a wide diversity of biological processes. These phase-separated droplets usually exhibit a viscoelastic mechanical response. A behavior rationalized by modeling the complex molecules that make up a condensate as stickers and spacers, which assemble into a network-like structure. The proper functioning of biocondensates requires precise control over their composition, size, and mechanical response. For example, several neurodegenerative diseases are associated with dysfunctional condensates that solidify over a long period of time (days) until they become solid. A phenomenon usually described as aging. The emergence of such a long timescale of evolution from microscopic events, as well as the associated microscopic reorganization leading to aging, remains mostly an open question. In this article, we explore the connection between the mechanical properties of the condensates and their microscopic structure. We propose a minimal model for the dynamic of stickers and spacers, and show that entropy minimization of spacers leads to an attractive force between stickers. Our system displays a surprisingly slow relaxation toward equilibrium, reminiscent of glassy systems and consistent with the liquid-to-solid transition observed. To explain this behavior, we study the clustering dynamic of stickers and successfully explain the origin of glassy relaxation.
Soft matter |||| Theoretical biophysics
preprint: APS/123-QED
Introduction
Liquid-liquid phase separation has emerged as a major organizing principle for the compartmentalization of cellular processes into membraneless organelles known as biomolecular condensates. Condensates are associated with various functions of the cell, although their precise role often remains elusive [1], to the point that it is not even clear if some of them are functional, incidental or pathological [2]. Similarly, little is known about the complete composition of condensates; however, one or several scaffolding components that drive the phase separation can usually be identified, including proteins , DNA [3, 4], or RNA [5, 6]. Despite their diversity, condensates also display common features. It is usually accepted that their formation requires weak multivalent interactions [7], mediated by disordered portions of proteins [8] or by RNA/DNA-protein interactions [9]. Similarly, micro-rheology experiments suggest that biocondensates exhibit a general viscoelastic mechanical response [10, 11]. This picture has been rationalized through the sticker and spacer model, which describes proteins as generic sequences of sticky (stickers) and non-sticky (spacers) regions [12]. The complex interactions that make up a condensate give rise to an emergent long-time transition to a solid-like state, usually described as aging [13, 14]. This process has been associated with protein denaturation [15], amyloid formation [16] and, from a physiological perspective, with neurodegenerative diseases [17]. Current physical pictures of condensate aging either rely on the modeling of specific proteins’ interaction – like $\beta$-sheet stacking [18] –, or do not model the microscopic dynamic of the components [19, 20]. Yet, aging seems to be a fairly general behavior, essentially insensitive to the details of the condensate’s composition, which suggests that a universal physical mechanism may govern this transition. A general microscopic picture explaining how large timescales (days or more) emerge from a collection of microscopic processes, that typically occurs on much shorter timescales, ($\mu s$) is missing.
In this work, we use the sticker and spacer framework to model the microscopic structural rearrangement taking place in condensates, and leading to aging. This approach essentially considers proteins as a heterogeneous collection of localized, attractive regions (stickers) and longer, neutral regions (spacers). Stickers reversibly associate via non-covalent interactions, while spacers provide flexibility and connection between stickers, resulting in the formation of a percolating network. The ensuing mechanical behavior of this network is typical of an associative gel, characterized by a viscoelastic response to stress. At short timescales, the gel displays solid, elastic-like behavior due to the entropic elasticity of the spacers, resisting deformation. Conversely, over longer timescales, the reversible binding and unbinding of stickers enable the material to flow, displaying liquid-like characteristics. Homogeneous associative gels display a soft relaxation over a single characteristic timescale, directly related to the unbinding dynamics of the stickers [21]. In contrast, heterogeneous gels feature multiple coexisting timescales, each corresponding to a different sub-region within the system, and a complex relaxation behavior emerges from their superposition [22]. Strong heterogeneity results in a broad distribution of timescales, sometime heavy tailed. In these extreme cases, a portion of these timescales are arbitrarely large and relaxation seemingly never ends, a phenomena typical of glasses, and known as aging [23, 24]. To investigate how aging can emerge from the collective rearrangement of the gel components, we model them as sticker and spacer. the former acting as diffusing particles that can bind to (and unbind from) the latter, that represent Gaussian polymers, possibly capturing intrinsically disordered regions of proteins or long RNA strands. Using a combination of analytical results and simulations, we show that an entropic Casimir-like force drives the clustering of stickers, in accordance with several other works [15, 3, 25]. We find that collective rearrangement of the stickers induce a slow-down of the dynamic of the system. As a consequence, the relaxation toward equilibrium is surprisingly slow and reminiscent of the usual glass dynamic [26]. To explain our observations, we introduce a minimal model for the dynamic of clusters, which provide a scaling law for the dynamical slowing-down. Despite its simplicity, our model displays a rich phenomenology leading to glassy dynamic which includes emergence of large timescales, ergodicity breaking and aging. By focusing on the minimal ingredients required for the emergence of a soft glassy relaxation, our model proposes a universal mechanism to aging of complex associative gels.
Results
Dynamic of stickers and spacers
We first introduce a physical model for the dynamic of stickers and spacers. Starting with spacers, which correspond to RNA or DNA strands, or to intrinsically disordered segments of proteins. Their dynamics are similar to those of a polymer, and we consider the dilute regime where excluded volume or topological constraints due to entanglement are negligible. Therefore, we model spacers as linear Gaussian polymers. In this work, we will focus on the dynamic of a single polymer strand and treat the others effectively. To model protein interactions, we consider that polymers interacts with a background disordered potential, effectively representing the surrounding environment as represented in Fig. 1. The positions of the minima of this potential correspond to the positions of the stickers. We propose that the long time evolution of the condensates results from the slow evolution of the background potential, that we model through the diffusion of stickers. While the diffusion dynamics of stickers may differ from a free particle depending on the exact nature of what they represent, we expect that the details of the microscopic dynamics will not significantly affect the long-time evolution of our system. Additionally, considering that stickers correspond to proteins or sticky regions of a protein, we treat them as effectively always bound to another polymer strand in the background. As a consequence, whenever the polymer binds to a sticker, it effectively creates a cross-link that impairs its diffusion and constraints the conformations of the polymer. Here, we consider the limiting case where stickers do not diffuse at all when bound to the polymer, thus creating an anchoring point. This dynamic provides enough freedom for the system to rearrange while capturing the constraining effect of binding.
Figure 1: Schematic representation of our sticker and spacer model. The disordered background potential is represented as a color coding, while stickers are represented as diffusing particles. We focus on a single polymer dynamic highlighted in bold, while the other effective polymers are in shaded. (a) Represents the initial state in which stickers are homogeneously spread in space. (b) Qualitatively represent a typical final state as predicted by our model.
To describe the dynamic of the stickers, we now consider the probability of the $i$-th sticker to be in position $\bm{r}_{i}$ at time $t$ in the bound state along the lineic position $\ell$ of the polymer denoted by $p_{\text{b}}(\bm{r}_{i},\ell,t)$. Similarly, we write the probability to be unbound $p_{\text{ub}}(\bm{r}_{i},t)$. Given the previously introduced dynamics, the time evolution of these two probability distributions is governed by the two coupled master equations:
$\partial_{t}p_{\text{b}}(\bm{r}_{i},\ell,t)=k_{\text{b}}(\bm{r}_{i},\ell)p_{\text{ub}}(\bm{r}_{i},t)-k_{ub}p_{\text{b}}(\bm{r}_{i},\ell,t)$ (1)
$\partial_{t}p_{\text{ub}}(\bm{r},t)=k_{ub}p_{\text{b}}(\bm{r}_{i},\ell,t)-k_{\text{b}}(\bm{r}_{i},\ell)p_{\text{ub}}(\bm{r}_{i},t)+D\nabla^{2}p_{ub}(\bm{r}_{i},t),$
where $k_{\text{b}}$ and $k_{\text{ub}}$ are respectively the binding and unbinding rates of the sticker with the considered polymer. We first write the unbinding rate using a single energy scale under the Kramers approximation:
$k_{\text{ub}}=\frac{1}{\tau_{0}}e^{-\beta E_{\text{b}}},$ (2)
where $\beta=1/(k_{B}T)$ is the inverse temperature, $E_{b}$ the binding energy scale, and $\tau_{0}$ a timescale associated with the microscopic escape process. While the complexity of real proteins’ interaction should lead to a distribution of binding energies, we show here that this assumption is not necessary for the system to display a glassy dynamic, and address the effect of a more complex model in the discussion section.
Concerning the binding rate, we assume that the disordered potential represented by the stickers is partially-quenched. Meaning that it evolves on a much larger timescale than the polymers. As a consequence, polymers have time to equilibrate between two binding/unbinding events. Assuming that all conformations of the polymer are equiprobable, we show in Supplementary Material that the probability for a given conformation of the polymer to meet a sticker located at $\bm{r}$ in a volume $a^{3}$ (where $a$ is the monomer size) at a lineic position $\ell$ is given by:
$p_{\text{meet}}(\bm{r},\ell)=e^{S_{\text{b}}(\bm{r},\ell)-S_{\text{ub}}}.$ (3)
where $S_{\text{b}}$ and $S_{\text{ub}}$ correspond to the entropy of the polymer strand when bound and unbound to the sticker, respectively. Our approximation of fast polymer conformational exploration means that the limiting timescale is the binding time. Thus, we write the binding rate of the sticker to the polymer as:
$k_{\text{b}}(\bm{r},\ell)=1/\tau_{0}\sum_{\alpha}p^{\alpha}_{\text{meet}}(\bm{r},\ell)=1/\tau_{0}\sum_{\alpha}e^{S^{\alpha}_{\text{b}}(\bm{r},\ell)-S^{\alpha}_{\text{ub}}},$ (4)
Where the sum over $\alpha$ represents the participation of all the polymer strands in the vicinity of the given sticker. Together with (2), this expression respects detailed balance, thus guaranteeing that the evolution of the system is consistent with thermodynamic equilibrium. Notice that $S^{\alpha}_{\text{b}}(\bm{r},\ell)$ introduces a coupling between the dynamic of the different stickers through $\ell$ that ultimately depends on where all the other stickers are bound (see Supplementary Material for the full expression). As a result, we cannot directly solve this system of equation without further approximation. In the next section, we use a Gillespie simulation to study the time evolution of our system.
Early time behavior
Figure 2: (a) Early time evolution of the Percentage of bound stickers in the system.Here, we use $L=2.10^{4}$ and $N=100$. (b) Plateau value of the percentage of bound stickers as a function of the binding energy, rescaled by the lineic density of stickers. The scattered points correspond to the measurements obtained from the panel (b).
We simulate a polymer of length $L$, which defines the system size, surrounded by $N$ initially unbound stickers spread homogeneously in space. First, looking at the early time evolution of the system, we observe in Fig. 2 (a) a rapid increase in the number of stickers bound to the polymer until a plateau is reached. We show in Fig. 2 (b) how the height of the plateau depends on the binding energy, observing a sharp increase from $0\%$ of the bound stickers to $100\%$ at a critical value of the binding energy. This phenomenon is reminiscent of the phase transition observed in the Poland-Scheraga model of DNA denaturation [27]. In their model, two DNA strands can bind to one another through sticky base pairs that are regularly spaced along the strands. Similarly to our model, the energetic binding energy must overcome the loss of entropy associated with the binding. As the linear density of stickers increases, it becomes increasingly favorable to bind them as the entropic cost for successive binding decreases. Their theory predicts a phase transition with a critical energy scaling, $E_{c}\propto\log(L/N)$. We observe that our model consistently displays a similar scaling based on the collapsed curves of Fig.2 As a consequence, for our system, low values of the binding energy (or high temperature) mean that stickers are mostly unbound, and the system mechanically responds like a liquid. If the value of the binding energy is higher than its critical value (or the temperature lower than its critical value), stickers remain bound most of the time, leading to a viscoelastic mechanical response similar to that of an associative gel. This type of gel displays a solid-like mechanical response at timescales lower than the unbinding time and liquid-like at timescales much larger than the unbinding time. In this work, we focus on the viscoelastic aging regime; for this reason, we always consider the binding energy above its critical value. In this case, unbinding becomes the time-limiting process for the dynamics of the system. In the following, we do not discuss the influence of the binding energy, which can be ruled out by rescaling times with respect to $1/k_{\text{ub}}$. In the high energy regime, the phenomena discussed are insensitive to the value of the binding energy.
Glassy relaxation
Figure 3: (a) sticker distance probability distribution at different time-point. The color coding corresponds to the time, and the analytical solution is provided as dotted line. We simulate systems with fixed lineic density: $N/L=5.10^{-2}$. Small systems of length $L=100$ successfully converge toward the equilibrium distribution. (b) Large systems, of length $L=2000$, evolves toward the analytic solution without reaching it. (c) We follow the equilibration process by looking at the decay of the free energy per unit length over time, the initial binding regime as been removed for improved visibility. Using a fixed lineic density of stickers, we expect all curves to converge to the same relative value. Large systems are stuck at a higher value of the free energy, stressing its inability to reach equilibrium. (d) Schematic representation of the partial equilibration of the system into sub-clusters.
Once all stickers are bound, the system’s energy becomes essentially fixed. However, the entropy of the polymer also depends on the positions of the stickers. We study the equilibration of the system through the diffusion of stickers. By setting the time derivative of the master equation to zero and independently equilibrating the binding/unbinding dynamics and the diffusion dynamics, we demonstrate in the Supplementary Material that the equilibrium probability distribution of finding two stickers interacting with a polymer of length $L$ at a distance $|r|$ from each other is given by:
$p(\bm{r})=\frac{3\text{erfc}\left(\sqrt{\frac{3}{2aL}}|\bm{r}|\right)}{2\pi aL|\bm{r}|}$ (5)
We derive an approximated formula in Supplementary Material for a system of $N$ particles, considering polymer strands that mediate interactions as a mean field. The simulated time evolution of the probability distribution toward the expected equilibrium analytic solution for smaller systems is depicted in Fig. 3 (a). Contrastingly, Fig. 3 (b) shows that larger systems, while converging similarly, do not reach the predicted equilibrium solution To further characterize the relaxation of our system toward equilibrium, we plot, in Fig. 3 (c), the free energy per unit length difference between our system and the theoretical free energy of an unconstrained polymer. Calculated as $f_{\text{free}}=F_{\text{free}}/L=NE_{b}/L-\log(4\pi)$ (where $4\pi$ is the solid angle between two monomers). We use a fixed lineic density of stickers ($N/L=5.10^{-2}$) and different system sizes, characterized by $L$. We observe a slow logarithmic decay of the free energy, followed by a plateau, where the total relaxation time depends on the system size. This type of logarithmic relaxation is usually associated with glassy systems [26], and we propose a model in the next section to better understand this unique behavior. Looking at the height of the plateau, we notice that it first increases with the system size before becoming independent of it. This is an unexpected behavior as the system is self-similar along the polymer, and the equilibrium free energy per unit length is an intensive quantity with respect to $L$. Additionally, we show in Supplementary Material that the height of the plateau also depends on the diffusion constant, $k_{\text{diff}}$. These results suggest that systems above a certain size are unable to fully relax to equilibrium. Instead, they form substructures, each equilibrated, that we call clusters. We propose in Fig. 3 (d) a 1D schematic of this picture. Starting from a homogeneous distribution of stickers, they aggregate into equilibrated clusters. If the total system size is similar to the typical cluster size, the plateau free energy is the equilibrium one. For a system much larger than the cluster size, the plateau free energy is higher in average and intensive again.
To further examine the formation of equilibrated subsystems, we employ a hierarchical clustering algorithm to group stickers into clusters. This algorithm uses a single maximum length to define connectivity, here, we use the characteristic lengthscale of Eq. (5) to match the definition of clusters as equilibrated subsystems. Fig. 4 (a) illustrates the time evolution of the relative cluster size, showing that smaller systems tend to converge into a single cluster, while larger systems form clusters of fixed size, effectively showing a decay in the relative size.
Origin of the glassy relaxation
We now investigate the origin of the slow logarithmic relaxation dynamics of the free energy. Logarithmic decay typically implies a continuous slowing down of the dynamics during relaxation. To confirm this, we examine the intermediate scattering function (ISF), defined as:
$I(\bm{k},t)=\sum_{j=0}^{N}\left<e^{i\bm{k}(\bm{r}_{i}(t)-\bm{r}_{j}(0)}\right>.$ (6)
The modulus of the ISF characterizes the spatial and temporal dependence of the relaxation dynamics, It is commonly used to study glassy relaxations [28]. Here, we focus on the wave vector associated with the inter-sticker distance,$k=1/<r>$, where $<r>$ is the average sticker distance computed from Eq. (5). In Fig. 4 (b), we plot the decay of the ISF over time, quantifying the decorrelation of the stickers’ positions after a lag time denoted $t_{\text{lag}}$ along the relaxation. Unlike an exponential decay consistent with a normal random walk, we observe a stretched exponential relaxation defined by $\exp((-t/\tau)^{\alpha})$, with $\alpha\approx 0.7$. This stretched exponential decay is characteristic of the $\alpha$ relaxation in glassy liquids [29, 30] and indicates a relaxation slower than exponential, usually associated with subdiffusive microscopic processes [31]. By fitting the measured ISF with a stretched exponential, we extract the characteristic relaxation time of the system. The inset of Fig. 4 (b) shows the increase of the characteristic relaxation time $\tau$ as a function of the simulation time, confirming our intuition of a dynamical slow-down.
We now propose a model for the origin of this dynamical slow-down. In the previous section, we demonstrated that system equilibration involves the growth of clusters. Individual exchanges of stickers between clusters do not modify the average size; to increase the average cluster size, it is necessary to dissolve another cluster. To compute the dissolution time of a cluster, we consider a system with $N$ stickers spread across clusters. The average number of stickers in a cluster is denoted $\bar{n}$. We model the exchange of stickers as a two-step process: unbinding from one cluster and then rebinding in a random cluster, selected randomly and with equal probability from all available clusters. Each exchange occurs according to a Poisson process with a constant rate $1/\tau_{\text{exch}}$, which effectively encapsulates a complex process of diffusion coupled to the binding/unbinding dynamic, that we do not investigate here. For a given cluster, the probability $P(n,t)$ to contain $n$ stickers at time $t$ evolves as:
$\frac{\text{d}P(n,t)}{\text{d}t}=$ $1/\tau_{\text{exch}}\left[P(n+1,t)(n+1)+P(n-1,t)\bar{n}\right.$ (7)
$\left.-P(n,t)(n+\bar{n})\right],$
derived from the master equation, where a cluster with $n$ stickers loses one at a rate $1/\tau_{\text{exch}}n$ and gains one from any of the $N/\bar{n}$ clusters at a rate $N/(N/\bar{n})/\tau_{\text{exch}}=\bar{n}/\tau_{\text{exch}}$. Assuming that the smallest cluster is the one dissolving first, and $N\rightarrow\infty$, we compute in the Supplementary Material the average dissolution time:
$\tau(\bar{n})\underset{N\rightarrow\infty}{\sim}\tau_{\text{exch}}\frac{e^{\bar{n}}}{\bar{n}^{2}},$ (8)
We now investigate the long-time collective growth of clusters resulting from the dissolution of other clusters. Assuming that clusters dissolve at a rate $1/\tau(\bar{n})$, and its sticker is redistributed among the $N/\bar{n}-1$ other clusters. We describe the dissolution events as continuous in time, and write a differential equation for the time evolution of the average cluster size:
$\frac{\text{d}\bar{n}}{\text{d}t}=1/\tau_{\text{exch}}e^{-\bar{n}}\frac{\bar{n}^{2}}{N/\bar{n}-1},$ (9)
This non-linear differential equation cannot be solved analytically; however, we show in the Supplementary Material that for a large systems ($N$ and $\bar{n}$ large) $\bar{n}\propto\log(t)$. This logarithmic scaling can be used together with the dominant factor of Eq. (8): $\tau(\bar{n})\propto e^{\bar{n}}$. Which gives $\tau(t)\propto t^{\gamma}$, where $\gamma$ is a constant that depends on $\tau_{\text{exch}}$. This power-law qualitatively agrees with the inset of Fig. 4 (b), which confirms the relation between the relaxation time of the ISF, and the dissolution time of growing clusters. Here, we focused on the dominant mathematical factors, and neglected many processes that could provide a more quantitative agreement with the simulation. This choice is motivated by the initial strong simplification of our sticker and spacer model. And we believe that only qualitative results will be robust to the addition of real microscopic details.
Figure 4: (a) average cluster size evolution, clusters are built using a hierarchical algorithm that use a minimum distance linkage criteria. The distance used is the average sticker distance computed using the pair correlation function of Eq. (5), in our case $\bar{d}=3.6$. (b) Modulus, of the intermediate scattering function for $|\bm{k}|=1/\bar{r}=0.37$, where $\bar{r}$ is the average distance between stickers. The stretched exponential fits are in good agreement, and the corresponding $\tau$ values obtained are plotted in the inset. The value of $\alpha$ does not evolve significantly over time, and remains $\approx 0.7$.
Discussion
We have introduced a minimal model to describe the dynamics of protein or RNA/DNA-protein condensates. By utilizing the general sticker and spacer framework, we examined the long-term structural evolution of these systems. Our investigation primarily focused on two general effects: the entropic cost associated with binding and the increasing difficulty of the system to rearrange upon equilibration. Our model reveals two distinct mechanisms: the emergence of a Casimir attractive force between stickers mediated by spacers and the clustering of stickers leading to glassy relaxation. The first mechanism is reminiscent of bridging-induced phase separation, where RNA or DNA-protein binding is driven by a similar entropic force, triggering condensation [3, 9]. In our framework, this entropic attractive force appears for high binding energy, as the simultaneous unbinding of multiple stickers would disrupt the attractive interaction This approximation might not be universally valid for all biological systems. in fact, we anticipate that in many cases, the Casimir force will not be the primary driver of structural reorganization. Nonetheless, in our model, this attractive force is only useful to drive sticker clustering, which is ultimately responsible for glassy relaxation. This clustering can also be driven by system-specific mechanisms, such as three-body stabilizing interactions or multiple $\beta$-sheet stacking, and has been observed in experimental systems [25, 15, 32]. In fact, such clustering is also a general behavior of random walkers in disordered media, commonly known as localization in this context [33]. For these reasons, we are confident that the clustering of stickers described here is a fairly universal feature, from which we have only highlighted the entropic minimization effect.
Our model shares similarities with previously studied models of random walker with trap [34, 35], which reproduce key features, such as glass transition, subdiffusion, or localization, reminiscent of the phenomenology observed here. However, these models use a prior distribution of binding energies (e.g., exponential), aging occurs through the exploration of deeper and deeper traps by the random walker. On the one hand, real proteins exhibit a distribution of interaction energy, which is sufficient to induce complex relaxation of a single protein [36]. On the other hand, proteins’ interactions –electrostatic, hydrophobic, or hydrogen bonding– are all of the same order of magnitude ($\approx 10k_{B}T$), which should lead to a fairly narrow distribution of interaction energy. For this reason, intrinsic protein complexity does not justify a priori the existence of ever-deepening traps needed for long-time aging. Our model offers a rationale for considering a broad distribution of energetic interactions, as proposed in previous studies [19, 37]. By focusing on the exponential part of Eq. (8), we find that the dissolution time of clusters is similar to a Kramer escape time for a trap of depth $\bar{n}$ time larger than an isolated sticker. Such mechanism is known for enhancing small heterogeneities of stickers, leading to complex viscoelastic relaxation [38].
Acknowledgements.
We thank Lara Koehler for her critical reading of the manuscript. This work was supported by the Swiss National Science Foundation under grant CRSII5 193740.

### v3

@@
-Microscopic model for aging of biocondensates
+Entropic Clustering of Stickers Induces Aging in Biocondensates
 Abstract
-Biomolecular condensates are membraneless compartments in the cell that are involved in a wide diversity of biological processes.
-These phase-separated droplets usually exhibit a viscoelastic mechanical response.
+Biomolecular condensates are cellular phase-separated droplets that usually exhibit a viscoelastic mechanical response.
 A behavior rationalized by modeling the complex molecules that make up a condensate as stickers and spacers, which assemble into a network-like structure.
-The proper functioning of biocondensates requires precise control over their composition, size, and mechanical response.
-For example, several neurodegenerative diseases are associated with dysfunctional condensates that solidify over a long period of time (days) until they become solid.
-A phenomenon usually described as aging.
-The emergence of such a long timescale of evolution from microscopic events, as well as the associated microscopic reorganization leading to aging, remains mostly an open question.
+Condensates usually exhibit a solidification over a long period of time (days), a phenomenon described as aging.
+The emergence of such a long timescale of evolution from microscopic processes, as well as the associated microscopic reorganization leading to aging, remains mostly an open question.
 In this article, we explore the connection between the mechanical properties of the condensates and their microscopic structure.
-We propose a minimal model for the dynamic of stickers and spacers, and show that entropy minimization of spacers leads to an attractive force between stickers.
+We propose a minimal model for the dynamic of stickers and spacers, and show that entropy maximization of spacers leads to an attractive force between stickers.
 Our system displays a surprisingly slow relaxation toward equilibrium, reminiscent of glassy systems and consistent with the liquid-to-solid transition observed.
 To explain this behavior, we study the clustering dynamic of stickers and successfully explain the origin of glassy relaxation.
-Soft matter |||| Theoretical biophysics
+Soft matter $|$ Theoretical biophysics
 preprint: APS/123-QED
@@
 Current physical pictures of condensate aging either rely on the modeling of specific proteins’ interaction – like $\beta$-sheet stacking [18] –, or do not model the microscopic dynamic of the components [19, 20].
-Yet, aging seems to be a fairly general behavior, essentially insensitive to the details of the condensate’s composition, which suggests that a universal physical mechanism may govern this transition.
+Yet, aging seems to be a fairly general behavior, suggesting that its occurrence does not depend on specific compositional details.
+Together with experimental evidences of a disordered-to-order transition [21] or the formation of substructures [15, 22] suggests that some of the features of this transition can be governed by a universal physical mechanism.
 A general microscopic picture explaining how large timescales (days or more) emerge from a collection of microscopic processes, that typically occurs on much shorter timescales, ($\mu s$) is missing.
 In this work, we use the sticker and spacer framework to model the microscopic structural rearrangement taking place in condensates, and leading to aging.
-This approach essentially considers proteins as a heterogeneous collection of localized, attractive regions (stickers) and longer, neutral regions (spacers).
+We do not investigate the phase separation process, instead, we consider the components of a droplet as a heterogeneous collection of localized, attractive regions (stickers) and longer, neutral regions (spacers).
 Stickers reversibly associate via non-covalent interactions, while spacers provide flexibility and connection between stickers, resulting in the formation of a percolating network.
 The ensuing mechanical behavior of this network is typical of an associative gel, characterized by a viscoelastic response to stress.
-At short timescales, the gel displays solid, elastic-like behavior due to the entropic elasticity of the spacers, resisting deformation.
-Conversely, over longer timescales, the reversible binding and unbinding of stickers enable the material to flow, displaying liquid-like characteristics.
+At timescales much smaller than the sticker lifetime, the gel displays solid, elastic-like behavior due to the entropic elasticity of the spacers, resisting deformation.
+Conversely, over longer timescales, the reversible binding and unbinding dynamic of stickers enable the material to flow, displaying liquid-like characteristics.
 Homogeneous associative gels display a soft relaxation over a single characteristic timescale, directly related to the unbinding dynamics of the stickers [23].
@@
 In these extreme cases, a portion of these timescales are arbitrarely large and relaxation seemingly never ends, a phenomena typical of glasses, and known as aging [25, 26].
-To investigate how aging can emerge from the collective rearrangement of the gel components, we model them as sticker and spacer.
+To investigate how aging can emerge from the collective rearrangement of the gel components, we model them as stickers and spacers.
 the former acting as diffusing particles that can bind to (and unbind from) the latter, that represent Gaussian polymers, possibly capturing intrinsically disordered regions of proteins or long RNA strands.
-Using a combination of analytical results and simulations, we show that an entropic Casimir-like force drives the clustering of stickers, in accordance with several other works [15, 3, 25].
+Using a combination of analytical results and simulations, we show that an entropic Casimir-like force drives the clustering of stickers, reminiscent of substructures experimentally observed [15, 3, 27].
 We find that collective rearrangement of the stickers induce a slow-down of the dynamic of the system.
@@
 To explain our observations, we introduce a minimal model for the dynamic of clusters, which provide a scaling law for the dynamical slowing-down.
+Finally, we compute the associated dynamic modulus of our system which demonstrate the long-time solidification resulting from aging.
 Despite its simplicity, our model displays a rich phenomenology leading to glassy dynamic which includes emergence of large timescales, ergodicity breaking and aging.
-By focusing on the minimal ingredients required for the emergence of a soft glassy relaxation, our model proposes a universal mechanism to aging of complex associative gels.
+By using a minimal model for the emergence of soft glassy relaxation, we propose a fairly general driving mechanism for the aging of biocondensates.
 Results
 Dynamic of stickers and spacers
-We first introduce a physical model for the dynamic of stickers and spacers.
+We consider a situation in which the system is already demixed, and focus on the dynamic of stickers and spacers.
 Starting with spacers, which correspond to RNA or DNA strands, or to intrinsically disordered segments of proteins.
-Their dynamics are similar to those of a polymer, and we consider the dilute regime where excluded volume or topological constraints due to entanglement are negligible.
+Their dynamics are similar to those of polymers, and we consider the dilute regime where excluded volume or topological constraints due to entanglement are negligible.
 Therefore, we model spacers as linear Gaussian polymers.
-In this work, we will focus on the dynamic of a single polymer strand and treat the others effectively.
-To model protein interactions, we consider that polymers interacts with a background disordered potential, effectively representing the surrounding environment as represented in Fig.
-1.
-The positions of the minima of this potential correspond to the positions of the stickers.
-We propose that the long time evolution of the condensates results from the slow evolution of the background potential, that we model through the diffusion of stickers.
-While the diffusion dynamics of stickers may differ from a free particle depending on the exact nature of what they represent, we expect that the details of the microscopic dynamics will not significantly affect the long-time evolution of our system.
-Additionally, considering that stickers correspond to proteins or sticky regions of a protein, we treat them as effectively always bound to another polymer strand in the background.
-As a consequence, whenever the polymer binds to a sticker, it effectively creates a cross-link that impairs its diffusion and constraints the conformations of the polymer.
-Here, we consider the limiting case where stickers do not diffuse at all when bound to the polymer, thus creating an anchoring point.
-This dynamic provides enough freedom for the system to rearrange while capturing the constraining effect of binding.
+Stickers are modeled as diffusive particles that can transiently bind to spacers, forming reversible cross-links.
+Stickers may represent adhesive domains within a polymer or separate binding proteins.
+We model both cases similarly by simulating a single polymer strand while treating the surrounding environment, composed of other polymers and stickers, as an effective bath of mobile particles that diffuse freely when unbound but become immobilized when bound to the simulated polymer.
+This dynamic for stickers reflects the formation of cross-links with other effective spacers that impair their diffusion and locally restrict polymer fluctuations as shown in Fig.
+1(a).
+Assuming that the dynamic of the one polymer simulated is representative of the others in the system, we are able to compute the viscoelastic response of such a polymer network.
+While the diffusion dynamics of stickers may differ from that of free particles depending on their exact nature, we will show that the long-time evolution of the system is controlled by the binding and unbinding dynamics.
+As a result, we expect that variations in the diffusion properties will not significantly affect our results.
+Here, we neglect the possibility that a sticker simultaneously interacts with other effectif spacers.
+This may lead to a slight overestimation of the number of free diffusing stickers or an underestimation of their effective diffusion.The goal of this modeling is to describe the essential physical trade-off: cross-linking lowers the system’s free energy but constrains its ability to rearrange
 Figure 1: Schematic representation of our sticker and spacer model.
@@
 (a) Represents the initial state in which stickers are homogeneously spread in space.
+The dynamics of stickers are also represented as small arrows: stickers bound to one polymer can diffuse or bind to another, whereas those bound to two polymers create a cross-link and must unbind before diffusing further.
 (b) Qualitatively represent a typical final state as predicted by our model.
-To describe the dynamic of the stickers, we now consider the probability of the $i$-th sticker to be in position $\bm{r}_{i}$ at time $t$ in the bound state along the lineic position $\ell$ of the polymer denoted by $p_{\text{b}}(\bm{r}_{i},\ell,t)$.
+To describe the dynamic of the stickers, we now consider the probability of the $i$-th sticker to be in position $\bm{r}_{i}$ at time $t$ in the bound state along the lineic position $\ell$ of the polymer considered denoted by $p_{\text{b}}(\bm{r}_{i},\ell,t)$.
 Similarly, we write the probability to be unbound $p_{\text{ub}}(\bm{r}_{i},t)$.
@@
 $\partial_{t}p_{\text{ub}}(\bm{r},t)=k_{ub}p_{\text{b}}(\bm{r}_{i},\ell,t)-k_{\text{b}}(\bm{r}_{i},\ell)p_{\text{ub}}(\bm{r}_{i},t)+D\nabla^{2}p_{ub}(\bm{r}_{i},t),$
-where $k_{\text{b}}$ and $k_{\text{ub}}$ are respectively the binding and unbinding rates of the sticker with the considered polymer.
+where $k_{\text{b}}$ and $k_{\text{ub}}$ are respectively the binding and unbinding rates of the sticker with the polymer.
 We first write the unbinding rate using a single energy scale under the Kramers approximation:
@@
 While the complexity of real proteins’ interaction should lead to a distribution of binding energies, we show here that this assumption is not necessary for the system to display a glassy dynamic, and address the effect of a more complex model in the discussion section.
-Concerning the binding rate, we assume that the disordered potential represented by the stickers is partially-quenched.
-Meaning that it evolves on a much larger timescale than the polymers.
-As a consequence, polymers have time to equilibrate between two binding/unbinding events.
-Assuming that all conformations of the polymer are equiprobable, we show in Supplementary Material that the probability for a given conformation of the polymer to meet a sticker located at $\bm{r}$ in a volume $a^{3}$ (where $a$ is the monomer size) at a lineic position $\ell$ is given by:
+Concerning the binding rate, we assumte that polymers have time to equilibrate between two binding/unbinding events.
+As a consequence, we represent the polymers through their equilibrium spatial probability distribution.
+We show in Supplementary Material that the probability for a given conformation of the polymer to meet a sticker located at $\bm{r}$ at a lineic position $\ell$ is given by:
 $p_{\text{meet}}(\bm{r},\ell)=e^{S_{\text{b}}(\bm{r},\ell)-S_{\text{ub}}}.$ (3)
@@
 Thus, we write the binding rate of the sticker to the polymer as:
-$k_{\text{b}}(\bm{r},\ell)=1/\tau_{0}\sum_{\alpha}p^{\alpha}_{\text{meet}}(\bm{r},\ell)=1/\tau_{0}\sum_{\alpha}e^{S^{\alpha}_{\text{b}}(\bm{r},\ell)-S^{\alpha}_{\text{ub}}},$ (4)
-Where the sum over $\alpha$ represents the participation of all the polymer strands in the vicinity of the given sticker.
+$k_{\text{b}}(\bm{r},\ell)=1/\tau_{0}p_{\text{meet}}(\bm{r},\ell)=1/\tau_{0}e^{S_{\text{b}}(\bm{r},\ell)-S_{\text{ub}}},$ (4)
 Together with (2), this expression respects detailed balance, thus guaranteeing that the evolution of the system is consistent with thermodynamic equilibrium.
-Notice that $S^{\alpha}_{\text{b}}(\bm{r},\ell)$ introduces a coupling between the dynamic of the different stickers through $\ell$ that ultimately depends on where all the other stickers are bound (see Supplementary Material for the full expression).
+Notice that $S_{\text{b}}(\bm{r},\ell)$ introduces a coupling between the dynamic of the different stickers through $\ell$ that ultimately depends on where all the other stickers are bound (see Supplementary Material for the full expression).
 As a result, we cannot directly solve this system of equation without further approximation.
@@
 Early time behavior
-Figure 2: (a) Early time evolution of the Percentage of bound stickers in the system.Here, we use $L=2.10^{4}$ and $N=100$.
+Figure 2: (a) Early time evolution of the Percentage of bound stickers in the system.
+Here, we use $L=2.10^{4}$ and $N=100$.
 (b) Plateau value of the percentage of bound stickers as a function of the binding energy, rescaled by the lineic density of stickers.
 The scattered points correspond to the measurements obtained from the panel (b).
-We simulate a polymer of length $L$, which defines the system size, surrounded by $N$ initially unbound stickers spread homogeneously in space.
-First, looking at the early time evolution of the system, we observe in Fig.
+We simulate a polymer of length $L$, which defines the system size, surrounded by $N$ initially unbound stickers spread homogeneously in a 3D space.
+Each Gillespie step consists in a move chosen among sticker diffusion, binding to a polymer in the vicinity of the sticker, or unbinding if it was bounded.
+All the rates are computed at the beginning of the step according to Eqs.
+(4) and (2), and a move is selected with a probability proportional to its rate.
+After selecting, and applying a specific move, a time increment is drawn from an exponential distribution with rate equal to the sum of rates of all the possible moves of the system.
+After each move, the rates are locally updated consistently with the new configuration of the system.
+Looking at the early time evolution of the system, we observe in Fig.
 2 (a) a rapid increase in the number of stickers bound to the polymer until a plateau is reached.
@@
 We simulate systems with fixed lineic density: $N/L=5.10^{-2}$.
-Small systems of length $L=100$ successfully converge toward the equilibrium distribution.
+Small systems of length $L=100$ successfully converge toward the predicted equilibrium distribution.
 (b) Large systems, of length $L=2000$, evolves toward the analytic solution without reaching it.
@@
 Using a fixed lineic density of stickers, we expect all curves to converge to the same relative value.
-Large systems are stuck at a higher value of the free energy, stressing its inability to reach equilibrium.
+However, large systems are stuck at a higher value of the free energy, stressing its inability to reach equilibrium.
 (d) Schematic representation of the partial equilibration of the system into sub-clusters.
@@
 We study the equilibration of the system through the diffusion of stickers.
-By setting the time derivative of the master equation to zero and independently equilibrating the binding/unbinding dynamics and the diffusion dynamics, we demonstrate in the Supplementary Material that the equilibrium probability distribution of finding two stickers interacting with a polymer of length $L$ at a distance $|r|$ from each other is given by:
-$p(\bm{r})=\frac{3\text{erfc}\left(\sqrt{\frac{3}{2aL}}|\bm{r}|\right)}{2\pi aL|\bm{r}|}$ (5)
-We derive an approximated formula in Supplementary Material for a system of $N$ particles, considering polymer strands that mediate interactions as a mean field.
+We first derive a mean field model of the equilibrium pair probability distribution for two stickers.
+To do so, we use the exact solution of the pair probability distribution between stickers derived in the Supplementary Material, and approximate that stickers share an equal portion of the polymer.
+We find that for a system with a density of stickers $L/N$, the probability of finding two stickers at a distance $|\bm{r}|$ from each other is given by:
+$p(\bm{r})=\frac{3\text{erfc}\left(\sqrt{\frac{3}{2aL/N}}|\bm{r}|\right)}{2\pi aL/N|\bm{r}|}$ (5)
 The simulated time evolution of the probability distribution toward the expected equilibrium analytic solution for smaller systems is depicted in Fig.
@@
 Contrastingly, Fig.
-3 (b) shows that larger systems, while converging similarly, do not reach the predicted equilibrium solution To further characterize the relaxation of our system toward equilibrium, we plot, in Fig.
-3 (c), the free energy per unit length difference between our system and the theoretical free energy of an unconstrained polymer.
+3 (b) shows that larger systems, while converging similarly, do not reach our approximated equilibrium solution.
+To determine whether the discrepancy arises from limitations in our mean field model or from the system’s inability to further equilibrate, we plot in Fig.
+3 (c) the free energy per unit length difference between our system and the theoretical free energy of an unconstrained polymer.
 Calculated as $f_{\text{free}}=F_{\text{free}}/L=NE_{b}/L-\log(4\pi)$ (where $4\pi$ is the solid angle between two monomers).
@@
 4 (a) illustrates the time evolution of the relative cluster size, showing that smaller systems tend to converge into a single cluster, while larger systems form clusters of fixed size, effectively showing a decay in the relative size.
+We show in Supplementary Material illustrative snapshots of the these systems.
 Origin of the glassy relaxation
@@
 The modulus of the ISF characterizes the spatial and temporal dependence of the relaxation dynamics, It is commonly used to study glassy relaxations [30].
-Here, we focus on the wave vector associated with the inter-sticker distance,$k=1/<r>$, where $<r>$ is the average sticker distance computed from Eq.
+Here, we focus on the wave vector associated with the inter-sticker distance,$k=1/<r>$, where <r><r> is the average sticker distance computed from Eq.
 (5).
@@
 We now propose a model for the origin of this dynamical slow-down.
-In the previous section, we demonstrated that system equilibration involves the growth of clusters.
+In the previous section, we showed that equilibration can be interpreted as the growth of clusters.
 Individual exchanges of stickers between clusters do not modify the average size; to increase the average cluster size, it is necessary to dissolve another cluster.
-To compute the dissolution time of a cluster, we consider a system with $N$ stickers spread across clusters.
+To compute the dissolution time of a cluster, we consider a system with $N$ stickers spread inside different clusters.
 The average number of stickers in a cluster is denoted $\bar{n}$.
-We model the exchange of stickers as a two-step process: unbinding from one cluster and then rebinding in a random cluster, selected randomly and with equal probability from all available clusters.
+We model the exchange of stickers as a two-step process: unbinding from one cluster and then rebinding in a random cluster, selected uniformly and with equal probability from all clusters.
 Each exchange occurs according to a Poisson process with a constant rate $1/\tau_{\text{exch}}$, which effectively encapsulates a complex process of diffusion coupled to the binding/unbinding dynamic, that we do not investigate here.
@@
 $\tau(\bar{n})\underset{N\rightarrow\infty}{\sim}\tau_{\text{exch}}\frac{e^{\bar{n}}}{\bar{n}^{2}},$ (8)
-We now investigate the long-time collective growth of clusters resulting from the dissolution of other clusters.
-Assuming that clusters dissolve at a rate $1/\tau(\bar{n})$, and its sticker is redistributed among the $N/\bar{n}-1$ other clusters.
+which shows a rapid, exponential growth of the dissolution time with the average cluster size that grows over time.
+To model the long-time collective growth of clusters resulting from the dissolution of other clusters, we assume that clusters dissolve at a rate $1/\tau(\bar{n})$, and its sticker is redistributed among the $N/\bar{n}-1$ other clusters.
 We describe the dissolution events as continuous in time, and write a differential equation for the time evolution of the average cluster size:
 $\frac{\text{d}\bar{n}}{\text{d}t}=1/\tau_{\text{exch}}e^{-\bar{n}}\frac{\bar{n}^{2}}{N/\bar{n}-1},$ (9)
-This non-linear differential equation cannot be solved analytically; however, we show in the Supplementary Material that for a large systems ($N$ and $\bar{n}$ large) $\bar{n}\propto\log(t)$.
+This non-linear differential equation cannot be solved analytically; however, we show in the Supplementary Material that for a large systems ($1\ll\bar{n}\ll N$ ) : $\bar{n}\propto\log(t)$.
 This logarithmic scaling can be used together with the dominant factor of Eq.
@@
 This choice is motivated by the initial strong simplification of our sticker and spacer model.
-And we believe that only qualitative results will be robust to the addition of real microscopic details.
+And we believe that only such qualitative results will be robust to the addition of real microscopic details.
 Figure 4: (a) Average cluster size evolution, clusters are built using a hierarchical algorithm that use a minimum distance linkage criteria.
@@
 The value of $\alpha$ does not evolve significantly over time, and remains $\approx 0.7$.
+Notice the abrut decay for small $t_{\text{lag}}$ that corresponds to the initial binding regime.
+Viscoelastic response
+We now use our measurements of the intermediate scattering function (ISF) to estimate the long-time evolution of the dynamic relaxation modulus of our system.
+The mechanical relaxation of a chemically crosslinked polymer gel occurs through the unbinding of stickers.
+When a loaded bond breaks, the stress stored in that cross-link is relieved.
+If the unbound sticker has sufficient mobility, future rebinding occurs without reintroducing stress into the network.
+In our model, bound stickers remain fixed, while unbound stickers diffuse.
+Consistent with our previous assumptions, we assume that unbound stickers diffuse rapidly compared to the relaxation dynamics of the polymer network.
+Based on these principles, we derive in the Supplementary Material a relation connecting the ISF to the dynamic modulus for wave vector $k=1/\langle r\rangle$:
+$G(\omega)=\frac{1}{\omega j}\int_{0}^{+\infty}e^{-j\omega t}I(t,t_{\text{lag}}),\text{d}t,$ (10)
+where $\Re[G(\omega)]=G^{\prime}(\omega)$ is the storage modulus, and $\Im[G(\omega)]=G^{\prime\prime}(\omega)$ is the loss modulus at angular frequency $\omega$.
+The results are shown in Fig.
+5 and are consistent with previous experimental [13] and simulation results [34].
+As expected from the growing relaxation timescales of the ISF, the crossover frequency—where the storage modulus exceeds the loss modulus, marking the transition from liquid-like to solid-like behavior-shifts to lower values as the lag time increases.
+This confirms that the material becomes progressively more solid-like at long times.
+To aid interpretation, we also provide reference scaling laws characteristic of Maxwell viscoelastic behavior.
+Although the measured curves resemble those of a Maxwell material, the scalings differ slightly.
+This deviation arises because the ISF follows a stretched exponential form rather than a single exponential.
+However, the stretching exponent $\alpha$ remains close to 1, leading to a viscoelastic response that closely approximates Maxwellian behavior.
+Figure 5: Estimated storage ($G^{\prime}$) and loss ($G^{\prime\prime}$) modulus, obtained from Eq.
+10, using ISF measured after different lag-time.
+Slopes characteristics of Maxwell viscoelastic gel are displayed as reference.
 Discussion
 We have introduced a minimal model to describe the dynamics of protein or RNA/DNA-protein condensates.
-By utilizing the general sticker and spacer framework, we examined the long-term structural evolution of these systems.
+By utilizing the general sticker and spacer framework, we examined the long-time structural evolution of these systems.
 Our investigation primarily focused on two general effects: the entropic cost associated with binding and the increasing difficulty of the system to rearrange upon equilibration.
@@
 The first mechanism is reminiscent of bridging-induced phase separation, where RNA or DNA-protein binding is driven by a similar entropic force, triggering condensation [3, 9].
-In our framework, this entropic attractive force appears for high binding energy, as the simultaneous unbinding of multiple stickers would disrupt the attractive interaction This approximation might not be universally valid for all biological systems.
+In our framework, this entropic attractive force appears for high binding energy, as the simultaneous unbinding of multiple stickers would disrupt the attractive interaction.
+This approximation might not be universally valid for all biological systems.
 in fact, we anticipate that in many cases, the Casimir force will not be the primary driver of structural reorganization.
@@
 On the one hand, real proteins exhibit a distribution of interaction energy, which is sufficient to induce complex relaxation of a single protein [38].
-On the other hand, proteins’ interactions –electrostatic, hydrophobic, or hydrogen bonding– are all of the same order of magnitude ($\approx 10k_{B}T$), which should lead to a fairly narrow distribution of interaction energy.
+On the other hand, proteins’ interactions, electrostatic, hydrophobic, or hydrogen bonding, are all of the same order of magnitude ($\approx 10k_{B}T$), which should lead to a fairly narrow distribution of interaction energy.
 For this reason, intrinsic protein complexity does not justify a priori the existence of ever-deepening traps needed for long-time aging.
@@
 Acknowledgements.
-We thank Lara Koehler for her critical reading of the manuscript. This work was supported by the Swiss National Science Foundation under grant CRSII5 193740.
+We thank Ned Wingreen for insightful discussions that contributed to improve this work and Lara Koehler for her thorough and critical reading of the manuscript.
+This research was supported by the Swiss National Science Foundation (grant CRSII5 193740).
+Data Availability
+All the data presented in this article come from c++ and python simulation.
+The source codes are available at : https://github.com/HugoLeRoy94/cpp_file_aging_condensate.git and https://github.com/HugoLeRoy94/Parallel_gillespie.git